# Jacobian Scopes: token-level causal attributions in LLMs

### v1

@@ Abstract
Large language models (LLMs) make next-token predictions based on clues present in their context, such as semantic descriptions and in-context examples. Yet, elucidating which prior tokens most strongly influence a given prediction remains challenging due to the proliferation of layers and attention heads in modern architectures. We propose Jacobian Scopes, a suite of gradient-based, token-level causal attribution methods for interpreting LLM predictions. By analyzing the linearized relations of final hidden state with respect to inputs, Jacobian Scopes quantify how input tokens influence a model’s prediction. We introduce three variants — Semantic, Fisher, and Temperature Scopes — which respectively target sensitivity of specific logits, the full predictive distribution, and model confidence (inverse temperature). Through case studies spanning instruction understanding, translation and in-context learning (ICL), we uncover interesting findings, such as when Jacobian Scopes point to implicit political biases. We believe that our proposed methods also shed light on recently debated mechanisms underlying in-context time-series forecasting. [^1] Our code and interactive demonstrations are publicly available at https://github.com/AntonioLiu97/JacobianScopes
@@ 1 Introduction
Large Language Models (LLMs) exhibit unexpected and remarkable emergent abilities, such as multi-step Huang and Chang (2023) and in-context learning (ICL) Dong et al. (2024); Garg et al. (2023). However, the mechanistic foundations of these emergent abilities currently remains a deep and open question Olsson et al. (2022); Reddy (2024); Von Oswald et al. (2023). As a step towards improving fundamental understanding in this regard, both industry and academia have been engaged in research aimed at interpreting LLMs reasoning process during such emergent thinking modes; as when solving Wei et al. (2022); Wang et al. (2023) or composing Wan et al. (2023); Guo et al. (2024). Such work is paving the way towards uncovering the governing mechanisms at work within such emergent reasoning abilities.
In this work, we present a suite of gradient-based causal attribution methods called Jacobian Scopes. Jacobian Scopes attribute a token prediction made by an LLM to key inputs. This is achieved by analyzing the Jacobian matrix which captures locally linearized relation between the predictive distribution with respect to a preceding input token vector.
We showcase the efficacy of our proposed Jacobian Scopes through a range of case studies, including natural language instruction such as system prompts and translation, as well as ICL tasks such as time-series continuation forecasting Liu et al. (2024); Gruver et al. (2023); Mirchandani et al. (2023). Across these examples, our proposed Jacobian Scopes reveal potential semantic reasoning processes, expose implicit political biases Bang et al. (2024), and hint at a potential mechanism for understanding how LLMs extrapolate stochastic and chaotic time-series data sets, in-context Zhang and Gilpin (2025).
@@ 2 Methodology
Notations. Let $\mathcal{H}:=\mathbb{R}^{d_{\text{model}}}$ denote the model’s hidden space. We represent a length-$T$ input sequence as $\bm{X}_{1:T}:=(\bm{x}_{1},\ldots,\bm{x}_{t},\ldots,\bm{x}_{T})$, where $\bm{x}_{t}\in\mathcal{H}$.
We view an auto-regressive LLM as a function that maps a sequence of inputs to the final-layer, post-layer-norm hidden state at the leading position:
@@
Forming the full Jacobian $\bm{J}_{t}$ requires $d_{\mathrm{model}}$ backward passes, which can be expensive for modern LLMs.
Fortunately, in practice we find that one often only needs to analyze the norm of $\bm{v}^{\intercal}\bm{J}_{t}$, where $\bm{v}\in\mathcal{H}$ captures a property of interest in the predictive distribution, such as confidence (effective inverse temperature) or logit of a specific word.
@@
Influence computation via auto-diff. For the leading prediction $\bm{y}$, the influence scores at all input positions $t\leq T$ can be computed with a single backward pass by auto-differentiating the customized scalar loss $\mathcal{L}(\bm{X}_{1:T}):=\bm{v}^{\intercal}\bm{y}$, as shown in Eq.˜2.
@@
In Section˜3, we present two specific choices for the direction of interest $\bm{v}$, leading to Semantic Scope and Temperature Scope. We also describe the more computationally intensive Fisher Scope, which instead requires the entire Jacobian $\bm{J}_{t}$.
3 Applications
This section discusses three formulations of Jacobian Scopes, each motivated by a specific attribution goal. All experiments are conducted on LlaMA-3.2 1B Grattafiori et al. (2024), unless otherwise specified.
@@ 3.1 Semantic Scope
which is exactly the predicted logit of the target token. Differentiating this loss using Eq.˜2 yields $\left\|\bm{w}_{\mathrm{target}}^{\intercal}\bm{J}_{t}\right\|_{2}$ as the influence score.
Intuitively, Semantic Scope quantifies how sensitive the target logit is to changes in specific input tokens.
Figs.˜1 and 2 demonstrate how Semantic Scopes may be employed to visualize semantic reasoning and implicit political bias in LLMs.
@@ 3.2 Fisher Scope
While Semantic Scope attributes the past tokens that affect a specific output logit, it does not capture their influence on the entire predictive distribution $\bm{p}(\cdot\mid\bm{X}_{1:T})$, which leads us to introduce Fisher Scope. The local information geometry around the predictive distribution $\bm{p}(\cdot\mid\bm{X}_{1:T})=\mathrm{softmax}(\bm{W}\bm{y})$ is characterized by the Fisher information matrix (FIM) Bishop and Nasrabadi (2006)
$\bm{F}_{u}=\bm{W}^{\intercal}\big(\mathrm{diag}(\bm{p})-\bm{pp}^{\intercal}\big)\bm{W},$
which defines a Riemannian metric that relates infinitesimal perturbations in $\bm{y}$ to second-order changes in the output distribution $\bm{p}$ Amari and Nagaoka (2000).
To assess how an input token $\bm{x}_{t}$ influences the model’s prediction, we then pull back Arvanitidis et al. (2022) this metric through the Jacobian $\bm{J}_{t}$ and obtain the FIM in input space:
$\bm{F}_{t}:=\bm{J}_{t}^{\intercal}\bm{F}_{u}\bm{J}_{t}.$ (4)
This FIM, detailed in Section˜A.7, defines a local quadratic form governing the change in predictive distribution induced by infinitesimal perturbations to the input Amari (2016):
$\mathrm{KL}\!\left(\bm{p}(\cdot\mid\bm{x}_{t}),\,\bm{p}(\cdot\mid\bm{x}_{t}+\bm{\delta})\right)\;\approx\;\frac{1}{2}\,\bm{\delta}^{\intercal}\bm{F}_{t}\bm{\delta}.$ (5)
In practice, we summarize this information by the trace of the Fisher matrix as
$\mathrm{Influence}_{t}:=\mathrm{tr}(\bm{F}_{t}),$ (6)
@@
Figure 3: Fisher Scope reveals that translated words, highlighted in red, are primarily influenced by their corresponding words in the source language shaded in blue.
As shown in Section˜A.8, $\mathrm{tr}(\bm{F}_{t})$ governs the expected KL divergence induced by an isotropic perturbation of the input. This scalar therefore summarizes how sensitive the model’s predictive distribution $\bm{p}$ is to input token $\bm{x}_{t}$.
Fig.˜3 displays the application of Fisher Scope on three translation tasks, revealing the word blocks in the source language that most affect the predictive distribution at target position $T$. Although Fisher Scope is well-motivated by information geometry and yields meaningful attribution in practice, it relies on the full Jacobian $\bm{J}_{t}$. Computing $\bm{J}_{t}$ requires one backward pass per hidden dimension, making Fisher Scope a factor of $d_{\mathrm{model}}$ more expensive than Semantic Scope. This could be prohibitive for long input contexts typical in ICL. In the next section, we use heuristics to derive a distribution-level sensitivity score that matches the computational cost of Semantic Scope, by exploiting the projection trick in Eq.˜1.
Table 1: Comparison of the Jacobian Scopes methods introduced.
Semantic Fisher Temperature
Quantity explained target logit $z$ predictive distribution $\bm{p}$ confidence $\|\bm{y}\|_{2}$
Influence score $\|\bm{w}_{\mathrm{target}}^{\intercal}\bm{J}_{t}\|_{2}$ $\mathrm{tr}(\bm{F}_{t})$ $\|\hat{\bm{y}}^{\intercal}\bm{J}_{t}\|_{2}$
Complexity (# forward passes) $O(1)$ $O(d_{\mathrm{model}})$ $O(1)$
@@ 3.3 Temperature Scope
Temperature Scope attributes which input tokens affect the model’s confidence, i.e., make the distribution more “peaked” in its support over the entire vocabulary $\mathcal{V}$.
Let the logit vector be $\bm{z}=\bm{W}\bm{y}\in\mathbb{R}^{|\mathcal{V}|}$, and subsequently decompose the hidden state into norm (scalar coefficient) and direction (unit vector), $\bm{y}=\|\bm{y}\|_{2}\hat{\bm{y}}$, where $\|\hat{\bm{y}}\|_{2}=1$. By linearity,
$\bm{z}=\bm{W}(\|\bm{y}\|_{2}\hat{\bm{y}})=\|\bm{y}\|_{2}\,(\bm{W}\hat{\bm{y}})=\beta_{\mathrm{eff}}\hat{\bm{z}}$,
@@
where we define the effective inverse temperature as $\beta_{\mathrm{eff}}:=\|\bm{y}\|_{2}$ and the normalized logits as $\hat{\bm{z}}:=\bm{W}\hat{\bm{y}}$.
The predicted distribution can thus be written as a Boltzmann distribution with effective inverse temperature $\beta_{\mathrm{eff}}$:
$p(y_{i}\mid\bm{X}_{1:T})=\mathrm{softmax}(\beta_{\mathrm{eff}}\hat{\bm{z}})_{i}.$
Increasing $\beta_{\mathrm{eff}}$ makes the distribution more peaked without changing the ranking of the logits given by $\hat{\bm{z}}$. Temperature Scopes uses the effective inverse temperature $\beta_{\mathrm{eff}}$ as the customized loss, $\mathcal{L}_{\mathrm{temperature}}=\beta_{\mathrm{eff}}=\|\bm{y}\|_{2},$ whose auto-differentiation yields the influence score:
$\mathrm{Influence}_{t}:=\left\|\hat{\bm{y}}^{\intercal}\bm{J}_{t}\right\|_{2}$ (7)
Similar to Fisher Scope, Temperature Scope measures the extent to which each input token influences the model’s entire predictive distribution. Thanks to its computational efficiency, Temperature Scope is particularly well-suited for attribution tasks involving long contexts, where the goal is to understand how an LLM forms its overall vocabulary predictions — such as when translating (see Fig.˜6 in Appendix) or extrapolating the behavior of time-series data Gruver et al. (2023).
@@
Figure 4: LLaMA-3.2 extrapolates a chaotic time-series via pattern matching. Temperature Scope reveals that when predicting future values in a partially observed Lorenz system, LLaMA-3.2 attends to regions in the input history that exhibit peaked patterns similar to those near the cutoff.
We investigate how LLaMA-3.2 extrapolates 1-dimensional stochastic and chaotic time-series trajectories, as studied in Liu et al. (2024); Zhang and Gilpin (2025); Gruver et al. (2023).
Given the observed history, which is a series of scalar states $\bm{s}_{1:T}=(s_{1},\ldots,s_{T}$), the next state $s_{T+1}$ in these trajectories follows a well-defined distribution.
Liu et al. (2024) showed that LLMs’ predicted next-state distributions $\bm{p}_{\mathrm{LLM}}(\cdot\mid\bm{s}_{1:T})$ steadily converges to the ground-truth transition rule $\bm{p}_{\mathrm{true}}(\cdot\mid\bm{s}_{1:T})$.
@@
Following Liu et al. (2024); Gruver et al. (2023), we tokenize the time-series data as alternating tokens of commas and 2-digit numbers, [^2] In LLaMA-3.2, there is one token for each 2-digit number from 10 to 99. yielding purely numerical prompts such as: $29,30,31,33,36,40,45,45,\ldots$ We then use Temperature Scope to visualize the influence of inputs (ignoring comma delimiters) on LLM’s predicted next-state distribution.
@@
As shown in Figs.˜4 and 8, when extrapolating chaotic time-series data with recurring but non-identical motifs, LLaMA exhibits consistent attribution patterns: it attends to segments in history that match the local pattern near the cutoff point. This attribution signature is in favor of the conjecture that LLMs might parrot the best-matching pattern in history via nearest-neighbor search in delayed embedding space Zhang and Gilpin (2025), which is a simple yet effective strategy that yields power-law scaling as observed by Liu et al. (2024).
Figure 5: Temperature Scope shows diminished influence of early context as LLaMA-3.2 extrapolates a Brownian motion.
@@
On the other hand, when faced with stochastic systems with no recurring motifs, such as the Brownian motion example shown in Figs.˜5 and 11, the LLM employs a different strategy: it preferentially attends to later parts of the context. This strategy explains the observation Liu et al. (2024) that the ICL loss curve plateaus early for unbounded stochastic systems such as Brownian motion and Geometric Brownian Motion: as such systems wander out of distribution, the LLM ceases recognizing earlier history as relevant, and gradually forgets earlier statistics. We investigate this hypothesis further in Section˜A.2.

### v2

@@ Abstract
Large language models (LLMs) make next-token predictions based on clues present in their context, such as semantic descriptions and in-context examples. Yet, elucidating which prior tokens most strongly influence a given prediction remains challenging due to the proliferation of layers and attention heads in modern architectures. We propose Jacobian Scopes, a suite of gradient-based, token-level causal attribution methods for interpreting LLM predictions. Grounded in perturbation theory and information geometry, Jacobian Scopes quantify how input tokens influence various aspects of a model’s prediction, such as specific logits, the full predictive distribution, and model uncertainty (effective temperature). Through case studies spanning instruction understanding, translation, and in-context learning (ICL), we demonstrate how Jacobian Scopes reveal implicit political biases, uncover word- and phrase-level translation strategies, and shed light on recently debated mechanisms underlying in-context time-series forecasting. To facilitate exploration of Jacobian Scopes on custom text, we open-source our implementations and provide a cloud-hosted interactive demo at https://huggingface.co/spaces/Typony/JacobianScopes.
@@ 1 Introduction
Large Language Models (LLMs) exhibit unexpected and remarkable emergent abilities, such as multi-step Huang and Chang (2023) and in-context learning (ICL) Dong et al. (2024); Garg et al. (2023). However, the mechanistic underpinning of these emergent abilities remains a deep and open question Olsson et al. (2022); Reddy (2024); Von Oswald et al. (2023).
This has fueled a growing endeavor, across academia and industry, to interpret and explain LLMs’ inference process during such emergent thinking modes; as when writing Wan et al. (2023); Guo et al. (2024) and reasoning Wei et al. (2022); Wang et al. (2023).
In this work, we present a suite of gradient-based causal attribution methods called Jacobian Scopes. Jacobian Scopes explains prediction made by an LLM by influential input tokens. This is achieved by analyzing the Jacobian matrix, which captures locally linearized relation between the predictive distribution and input embeddings.
Jacobian Scopes come in three variants: Semantic, Fisher, and Temperature Scopes, each corresponding to a different explanandum (attribution objective): logit of a certain token, the full predictive distribution, and the model’s uncertainty (effective temperature). A survey of related gradient-based attribution methods and their relation to Jacobian Scopes is provided in Section˜A.2.
We showcase the efficacy of our proposed Jacobian Scopes through a range of case studies, including natural language instruction such as system prompts and translation, as well as ICL tasks such as time-series forecasting Gruver et al. (2023); Mirchandani et al. (2023). We further evaluate attribution faithfulness quantitatively on two benchmark datasets (LAMBADA and IWSLT2017) across two suites of leading LLMs: LLaMA-3.2 Grattafiori et al. (2024) and Qwen2.5 Qwen Team (2025).
@@ 2 Methodology
Notations. We use bold lowercase letters (e.g. $\bm{x}$) for vectors and bold uppercase letters (e.g. $\bm{W}$) for matrices.
Let $\mathcal{H}:=\mathbb{R}^{d_{\text{model}}}$ denote the model’s hidden space. We represent a length-$T$ input sequence as $\bm{X}_{1:T}:=(\bm{x}_{1},\ldots,\bm{x}_{t},\ldots,\bm{x}_{T})$, where $\bm{x}_{t}\in\mathcal{H}$. We view an auto-regressive LLM as a function that maps a sequence of inputs to the final-layer, post-layer-norm hidden state at the leading position:
@@
Forming the full Jacobian $\bm{J}_{t}$ requires $d_{\mathrm{model}}$ backward passes, which can be expensive for modern LLMs. Fortunately, in practice we find it sufficient to compute the vector-Jacobian product, $\bm{v}^{\intercal}\bm{J}_{t}$, where $\bm{v}\in\mathcal{H}$ captures a property of interest in the predictive distribution, such as confidence (effective inverse temperature) or the embedding of a target token.
@@
Influence computation via auto-diff. For the leading prediction $\bm{y}$, the influence scores at all input positions $t\leq T$ can be computed with a single backward pass by auto-differentiating the customized scalar loss $\mathcal{L}(\bm{X}_{1:T}):=\bm{v}^{\intercal}\bm{y}$:
@@
In Section˜3, we present three specific choices for the direction of interest $\bm{v}$, leading to Semantic Scope, Fisher Scope and Temperature Scope.
3 Example applications
This section discusses three formulations of Jacobian Scopes, each motivated by a specific attribution objective. All experiments are conducted on LLaMA-3.2 1B, unless otherwise specified.
@@ 3.1 Semantic Scope
which is exactly the predicted logit of the target token. Differentiating this loss using Eq.˜2 yields the Semantic Scope influence score
$\mathrm{Influence}_{t}^{\mathrm{Sem}}:=\left\|\bm{w}_{\mathrm{target}}^{\intercal}\bm{J}_{t}\right\|_{2}.$ (4)
Intuitively, Semantic Scope quantifies how sensitive the target logit is to changes in specific input tokens. Figs.˜1 and 2 demonstrate how Semantic Scopes may be employed to visualize semantic reasoning and implicit political bias in LLMs.
@@ 3.2 Fisher Scope
Semantic Scope attributes the past tokens that affect a specific output logit, but it does not capture their influence on the entire predictive distribution $\bm{p}(\cdot\mid\bm{X}_{1:T})$. The latter motivates Fisher Scope, which is particularly suited for attribution tasks where the correct next-word prediction is non-unique, such as translation (see Fig.˜6(b)).
Fisher Scope is inspired by information geometry Fisher (1922), which studies how infinitesimal changes in model parameters affect the predictive distributions Bishop and Nasrabadi (2006). In our setting, the hidden state $\bm{y}$ parameterizes the family $\bm{p}(\bm{y})=\mathrm{softmax}(\bm{W}\bm{y})$, and the FIM $\bm{F}$, derived in Section˜A.9, measures how sensitively $\bm{p}$ responds to perturbations in $\bm{y}$:
$\bm{F}=\bm{W}^{\intercal}\big(\mathrm{diag}(\bm{p})-\bm{pp}^{\intercal}\big)\bm{W}.$
Intuitively, the spectrum of $\bm{F}$ reveals how sensitively a perturbation along each eigendirection can change the predictive distribution. Formally, the eigen-decomposition
$\bm{F}=\bm{U}\bm{\Lambda}\bm{U}^{\intercal},$
assigns to each direction $\bm{u}_{i}$ a scalar sensitivity $\lambda_{i}$, quantifying how strongly a perturbation along $\bm{u}_{i}$ changes $\bm{p}(\cdot\mid\bm{X}_{1:T})$, as measured by KL divergence Amari and Nagaoka (2000).
We call the leading eigenvector $\bm{u}_{1}$ the principal Fisher direction: it is the single direction in output space $\mathcal{V}$ along which a perturbation to $\bm{y}$ most sensitively changes $p(\cdot\mid X_{1:T})$. The Fisher Scope influence score is then defined as the projection of the Jacobian onto this direction:
$\mathrm{Influence}_{t}^{\mathrm{Fisher}}:=\|\bm{u}_{1}^{\top}\bm{J}_{t}\|_{2}.$ (5)
In Section˜A.10, we show that Fisher Scope can be viewed as a rank-1 approximation of the total information between predictive distribution $\bm{p}$ and input embedding $\bm{x}_{t}$.
@@
Figure 3: Fisher Scope reveals that translated words, highlighted in red, are primarily influenced by their lexical counterparts in the source language.
Figs.˜3 and 6 demonstrate how Fisher Scope reveals the translation strategies employed by LLaMA-3.2: Fig.˜3 illustrates word-level translation, where prediction at each location is primarily influenced by its direct lexical counterpart in the source language. We refer the reader to Section˜A.3 for additional examples where phrase-level translation requires reasoning across multiple source tokens.
Table 1: Comparison of the Jacobian Scopes methods introduced. $\bm{w}_{\mathrm{target}}$ denotes the target vocab embedding, $\bm{u}_{1}$ the principal Fisher direction, and $\hat{\bm{y}}$ the normalized hidden direction.
Semantic Scope Fisher Scope Temperature Scope
Quantity explained target logit $z$ full distribution $\bm{p}$ confidence $\|\bm{y}\|_{2}$
Influence score $\|\bm{w}_{\mathrm{target}}^{\intercal}\bm{J}_{t}\|_{2}$ $\|\bm{u}_{1}^{\intercal}\bm{J}_{t}\|_{2}$ $\|\hat{\bm{y}}^{\intercal}\bm{J}_{t}\|_{2}$
Recommended use case semantic reasoning & bias analysis NLP tasks with non-unique predictions in-context learning on numerical data
@@ 3.3 Temperature Scope
While semantic next-word predictions span a large and heterogeneous vocabulary, structured ICL tasks, such as time-series prediction, Gruver et al. (2023) often exhibit striking regularities: the predicted distributions form near-Gaussian peaks over a small subset of numerical tokens (see Figs.˜4 and 12 as well as Liu et al. (2025)). Temperature Scope identifies which input tokens control the width of this Gaussian peak, and hence the model’s forecasting uncertainty.
To define Temperature Scope, we first decompose the hidden state $\bm{y}$ into its norm and direction, $\bm{y}=\|\bm{y}\|_{2}\hat{\bm{y}}$, so that the logit vector could be written as
$\bm{z}=\bm{W}\bm{y}=\|\bm{y}\|_{2}\,(\bm{W}\hat{\bm{y}})=\beta_{\mathrm{eff}}\,\hat{\bm{z}}$,
@@
where $\beta_{\mathrm{eff}}:=\|\bm{y}\|_{2}$ is the effective inverse temperature and $\hat{\bm{z}}$ the normalized logits.
[…]
$\mathrm{Influence}_{t}^{\mathrm{Temp}}:=\left\|\hat{\bm{y}}^{\intercal}\bm{J}_{t}\right\|_{2}$ (6)
@@
Figure 4: Left: Temperature Scope reveals different strategies employed by LLaMA-3.2 during time-series forecasting: (a) when continuing the dynamics of a partially observed Lorenz system, LLaMA attends to regions in the input history that exhibit patterns similar to those near the cutoff. (b) When extrapolating unstructured Brownian motion, LLaMA shows diminished influence of early context. Right: Predicted probability distribution over the top 30 most probable tokens.
@@
To probe how an LLM reasons over its inputs during numerical ICL tasks, we follow Gruver et al. (2023) and tokenize time-series data as alternating commas and 2-digit numbers, [^1] In LLaMA-3.2, there is one token for each 2-digit number from 10 to 99. yielding purely numerical prompts such as $29,30,45,\ldots$ Temperature Scope then attributes the confidence of the predicted next-state distribution to the inputs.
@@
As shown in Fig.˜4, when extrapolating chaotic time-series data with recurring but non-identical motifs, LLaMA exhibits consistent attribution patterns: it attends to segments in history that match the local pattern near the cutoff point. This attribution signature directly supports the recent conjecture that LLMs might parrot the best-matching pattern in history via nearest-neighbor search in delayed embedding space Zhang and Gilpin (2025).
@@
On the other hand, when faced with stochastic systems with no recurring motifs, such as Brownian motion, the LLM employs a different strategy: it preferentially attends to later parts of the context. This strategy explains the observation Liu et al. (2024) that the ICL loss curve plateaus early for unbounded stochastic systems such as Brownian motion and Geometric Brownian Motion: as such systems wander out of distribution, the LLM ceases recognizing earlier history as relevant, and gradually forgets earlier statistics. We investigate this hypothesis further in Section˜A.4.
Section˜A.5 further shows that Temperature Scope produces more interpretable attributions of in-context forecasting behaviour than Semantic and Fisher Scopes.
4 Quantitative Evaluations
To complement the qualitative case studies in Section˜3, we provide quantitative evaluations against established attribution baselines using the area over the perturbation curve (AOPC) Samek et al. (2016), which measures the drop in the ground-truth target token’s log-probability when the top-kk% most influential tokens are ablated to zero vectors. A more negative AUPC indicates higher attribution quality.
[…]
Table˜2 reports AOPC for LLaMA-3.2 3B. Jacobian Scopes consistently outperform random ablation and Integrated Gradients across both datasets, and are on par with Input $\times$ Gradient — a pattern that holds consistently across all four models tested (see Tables˜3, 4 and 5).
We note that the goal of this evaluation is not to claim state-of-the-art performance on the AOPC score, but to demonstrate that Jacobian Scopes produce attributions that match widely adopted baselines across diverse tasks. The primary strength of Jacobian Scopes resides in firm mathematical grounding, geometric interpretability, and straightforward single-backward-pass implementation.
It is remarkable that AOPC — a counterfactual, interventional metric — validates locally-computed influence scores. This suggests first-order sensitivity as a reliable proxy for token-level causal relevance.
[…]
Method LAMBADA IWSLT2017
Random $0.23\pm 0.01$ $0.19\pm 0.01$
Integrated Grads $0.67\pm 0.01$ $0.58\pm 0.01$
Input $\times$ Grad $1.12\pm 0.01$ $0.77\pm 0.01$
Semantic Scope $1.16\pm 0.01$ $0.78\pm 0.01$
Temperature Scope $\mathbf{-1.17\pm 0.01}$ $0.76\pm 0.01$
Fisher Scope $\mathbf{-1.17\pm 0.01}$ $\mathbf{-0.80\pm 0.01}$